OpenStudy (anonymous):

A hotel is offering two weekend specials described below. plan 1: 3 nights and 4 meals for $233 Plan 2: 3 nights and 3 meals for $226.50 For accounting purposes, the hotel will record income for the stay and meals separately. so the cost per night for each special must be the same, and the cost per meal of each special must be the same. a. What is a system of equation for the equation? B. solve the system. What is the cost per night and the cost per meal?

OpenStudy (goformit100):

For Solving this Question, I would like to ask you that, do you have the detailed conception of : Profit and Loss

OpenStudy (anonymous):

No

OpenStudy (texaschic101):

3n + 4m = 233 -- >(-1) 3n + 4m = 233 3n + 3m = 226.50 ----------------- -3n - 4m = -233 (result of multiplying by -1) 3n + 3m = 226.50 ----------------add -m = -6.50 m = 6.50 now sub 6.50 in for m in either equation... 3n + 4m = 233 3n + 4(6.50) = 233 3n + 26 = 233 3n = 233 - 26 3n = 207 n = 207/3 n = 69 check... 3n + 3m = 226.5 3(69) + 3(6.5) = 226.50 207 + 19.50 = 226.50 226.50 = 226.50 (correct) ANSWER : cost per night (n) = $ 69.00 cost per meal (m) = $ 6.50

OpenStudy (texaschic101):

I just set up the equations and solved by using process of elimination :) Do you have any questions ?

OpenStudy (anonymous):

can you help me solve this problem?

OpenStudy (texaschic101):

What problem ? I solved the one you just asked. You have another one ?